OpenStudy (anonymous):

6-2(x-y)= 3x + 5 a. 5x -2y =-1 b 5x + 2y = 1 c 2x - 5y = 1 d 5x -2y = 1

OpenStudy (anonymous):

Standard form??

OpenStudy (nurali):

6-2(x-y)= 3x + 5 6-2x+2y=3x+5 6-2x-3x+2y-5=0 1-5x+2y=0 5x-2y=1